OpenStudy (anonymous):

f(x)= (7x^2+5x+4)/√x find f'(x). find f'(4). im not quite sure im doing this problem correctly

OpenStudy (anonymous):

What is the quotient rule? How can we apply it to f'(x)?

OpenStudy (anonymous):

f'g+fg"/g^2

OpenStudy (anonymous):

would i use the product rule for the top? and then the quotient rule?

OpenStudy (anonymous):

\[f(x) = \frac{7x^2 + 5x + 4}{x^{1/2}} \\ f'(x) = \frac{(14x + 5)(x^{1/2}) + (7x^2 + 5x + 4)(\frac{1}{2} x^{-1/2})}{x}\] Evaluate.

OpenStudy (anonymous):

thanks @bloopman